Two jets are traveling toward each other and are 4000 km apart. The rate of one jet is 100 km/h faster than the rate of the other. If the jets pass each other after 2.5 hours, what is the rate of the faster jet?
Answer by robertb(4012)

2.5x + 2.5(x+ 100) = 4000
===> 2.5x + 2.5x +250 = 4000
===> 5x = 3750
===> x = 750 kph, the rate of one jet,
x + 100 = 850 kph, the rate of the faster jet.